# Generalizable Graph Neural Networks for Robust Power Grid Topology Control

### v2

@@ 3.4 Regimes
We consider three environmental regimes, which represent difficulty levels for grid operation. The full-network regime involves the whole network. Lines are statically disabled in the planned-outage regime. The unplanned-outage regime introduces an opponent that disables lines spontaneously. We use the opponent specified by Manczak et al. [24]. The opponent disables a randomly selected line for four hours twice a day. A cooldown period ensures that outages are separated by at least an hour.
@@ 4.1 Datasets
We use the data generated with the two expert agents from our previous study [15]. The first agent is the greedy agent, which decreases the line loading greedily. The second agent is the N-1 agent, which pursues topologies that are N-1 redundant. We found that the N-1 agent performed superiorly to the Greedy agent in settings with and without outages. Both agents have an activity threshold, below which do-nothing actions are taken automatically. State-action pairs with such do-nothing actions or from unsuccessful days are excluded. Do-nothing actions above this threshold are included.
We create two datasets. First, a in-distribution (ID) dataset containing networks on which the models are trained and evaluated. This dataset was used in our previous study [15]. This dataset combines the data from the greedy agent and the N-1 agent. The data from the N-1 agent on the full network is used. The data from the greedy agent on a subset of the N-1 networks is used. These are the N-1 networks with lines 0, 2, 4, 5, 6, and 12 disabled, which the greedy agent can operate well [15].
@@
On both datasets, we split the datapoints into 70/10/20 train/validation/test sets based on their scenario. The sizes of the partitions are listed in Table 3. In this paper, ID or OOD networks refer to the networks in either the ID or OOD datasets. ID or OOD outages refer to the outages present in either ID or OOD networks.
@@
Table 3: The number of datapoints in the datasets and train/val/test sets.
@@ 4.2 Datapoints
Each datapoint includes the features per object, the topology vector, and the expert action. The features for each object type are listed in Table 4. The object features are normalized. If a line is disabled, its endpoint features are zero-imputed, and the corresponding topology vector values are -1. Actions are converted from a set-action format into a switch-action format. Thus, actions are represented by a vector with a length of the topology vector. Each index indicates whether the corresponding object is switched between busbars (a value of 1) or not (a value of 0). This presents a multi-label binary classification task.
@@ 4.3 FCNN
The fully connected neural network (FCNN) consists of an input layer, multiple hidden layers, and an output layer. The input vector is obtained by flattening the input features into a vector and concatenating the topology vector. The hidden layers use the ReLU activation function. The output layer uses the sigmoid activation function to constrain the output to the $(0,1)$ range. The output vector has the length of the topology vector, so one value $p_{u}\in(0,1)$ is predicted per node $u$.
@@ 4.4 Homogeneous GNN
Applying a graph neural network to the power grid requires a graph representation of the power grid. We represent the grid objects as nodes. We connect nodes if the corresponding objects are attached to the same busbar or are corresponding line endpoints. We call this graph representation homogeneous as it does not consider edge types. We call the associated GNN the HomGNN. Figures 2a and 2b display an example grid and its homogeneous representation.
@@
$\textbf{h}_{u,0}=\begin{cases}\text{MLP}_{gen}(\textbf{x}_{u})&\text{node $u$ represents a generator}\\ \text{MLP}_{load}(\textbf{x}_{u})&\text{node $u$ represents a load}\\ \text{MLP}_{line}(\textbf{x}_{u})&\text{node $u$ represents a line endpoint},\end{cases}$
where $\textbf{h}_{u,0}$ and $\textbf{x}_{u}$ are, respectively, the initial node embedding and the features of node $u$. The embeddings of subsequent layers are computed with the message passing rule [36]:
@@ 4.5 Busbar Information Asymmetry
The aforementioned graph neural network formulation has a problem. Each node’s output specifies whether to switch the corresponding object between busbars. However, messages from objects on the current busbar are passed, while messages from the objects on the other busbar are not. We call this the busbar information asymmetry. This is intuitively problematic as a switching decision requires information about the other busbar. This also decreases the model’s expressiveness. Objects at the same substation but on different busbars and objects at entirely different substations cannot be distinguished. Moreover, the absence of inter-busbar connections also decreases the graph’s connectivity. This can result in longer paths that may require deeper GNNs.
@@ 4.6 Heterogeneous GNN
Our solution is to represent the grid with a heterogeneous graph representation. The associated GNN is called the HetGNN. This heterogeneous representation features three edge types: one for objects on the same busbar, one for objects on the other busbar (but at the same substation), and one for corresponding line endpoints. This is displayed in Figure 2c. The message passing rule is adapted to consider a weight and neighborhood aggregation per edge type:
@@
where $\mathcal{N}_{same}$ and $\mathcal{N}_{other}$ denote the neighborhoods of objects on the same and other busbar, respectively. The $W_{line,k}\textbf{h}_{line,k}$ term represents the addition of the embedding of a node connected by a power line. It is therefore omitted for nodes that do not represent line endpoints. Figure 2e displays an example of heterogeneous message passing.
@@ 4.7 Optimization & Postprocessing
$_{i}=\begin{cases}1&\text{object $i$ corresponds to the target substation}\\ 1&\text{object $i$ corresponds to the predicted substation}\\ \alpha&\text{otherwise}\end{cases}\quad\forall w_{i}\in\textbf{w}.$ w .
There is no target substation if the target action is a do-nothing action. The predicted action is classified as a do-nothing action, without a predicted substation, if all predictions $_{i}\in\textbf{p}$ p do not exceed $0.5$. Otherwise, the predicted substation is the substation where the predictions $\textbf{p}_{s}$ at that substation $s\in S$ maximize
@@ 4.8 Hyperparameter Tuning & Training
Figure 3: The training curves of the five models per model type. Green lines show the HetGNNs, yellow lines the FCNNs, and blue lines the HomGNNs.
@@ 4.9 Imitation Learning Agents
The ML models are applied to the environment. They can be applied directly or combined with simulation. The naive agent executes a model’s predicted action directly. We observe that this agent occasionally fails by predicting a singular erroneous action. The verify agent addresses this by verifying predicted actions with simulation. The predicted action is simulated. A do-nothing action is selected if the simulated line loadings are increased beyond the thermal limit. We also consider hybrid agents. The verify+greedy agent normally functions as the verify agent, but switches to the greedy agent if a line breaches the thermal limit. The verify+N-1 agent functions similarly but with the N-1 agent. Each agent also uses the activity threshold used by the expert agents [15].
@@ 5.1 Supervised Learning
Table 6 lists the accuracies of the different models on the various data partitions and the ID/OOD network groups. Figure 4 shows the test accuracies of the models on both datasets. On the ID dataset (rows 1 to 3), the accuracies remain limited to approximately 80%. On each split, the HetGNN achieves the highest accuracy, followed by the FCNN and, lastly, the HomGNN. Each model type shows a tiny drop in accuracy between the train, validation, and test sets, indicating slight overfitting. Row 4 displays that accuracy drops slightly without the postprocessing step described in Sec. 4.7.
The default topology appears frequently in the ID dataset, comprising 46% of the test set. In this topology, all objects are attached to the same busbar. This avoids the busbar asymmetry problem and consequently negates the theoretical advantage of the HetGNN over the HomGNN. As shown in row 5, the accuracies on the standard topology are relatively similar. As shown in row 6, the accuracies vary considerably between topologies with split busbars.
@@
The GNNs obtain far higher accuracies on the OOD dataset than the FCNNs (rows 7 to 9). The HetGNNs also obtain higher accuracies than the HomGNNs. However, the performance of either the HomGNN or the HetGNN on the OOD dataset is still substantially lower than models trained on OOD data (column ’OOD-GNN’).
@@ 5.2 Error Analysis
We repeated the error analysis previously performed on the FCNN [15] to the two types of GNN models. We again find that the class imbalance and class overlap play a role in the limited accuracy. Figure 5 shows that both GNN types predict infrequent classes disproportionately infrequently. We investigated pairs of classes that were often confused, i.e., pairs of classes that were often mistaken for one another in prediction. Figure 6 shows that the errors of frequently confused classes are in overlapping regions.
@@
Inspection of the nearest neighbors supports the notion that the class overlap contributes to low accuracy. We applied one of the HetGNNs to a subset of 2,500 validation data points from the full network. The accuracy over the data points whose nearest neighbor is in the same class is 92.97%, but only 44.44% for those whose nearest neighbor is in another class. Other models showed similar results.
@@ 5.3 Graph Smoothness
The tendency of graph neural networks to suffer from oversmoothing is well-documented [3, 23]. We compute the mean average distance (MAD) values to measure graph smoothness [3]. MAD values quantify the similarity of node embeddings. Small MAD values indicate that node embeddings are similar and, thus, that the graph is smooth.
The MAD values are computed between neighbor nodes over the first 1000 validation datapoints. Figure 5.4 shows the MAD values between the various layers. As visible, the MAD values in the middle GNN layers are consistently higher in the HetGNN than in the HomGNN. This suggests that the heterogeneous representation leads to a less smooth local embedding. This finding is consistent among the models.
@@ 5.4 Graph Diameter
We mentioned that the heterogeneous graph representation can lead to shorter paths. We measure this by the diameter, i.e., the shortest path length between the two most distant nodes. We calculate the diameter by finding the lowest exponent of the adjacency matrix that produces a matrix without zero entries [7]. The diameters of the 25 most common topologies are shown in Figure 5.4. As visible, the heterogeneous graph representation infrequently leads to a shorter diameter.
@@ 5.5 Simulation Performance
Table 7 shows the days completed by the various agents in the different settings. The ’Full Network’ column describes the setting without outages. The ’ID Outages’ describes a setting with an opponent that causes random outages present in the ID dataset (see Sec. 4.1). The ’OOD Outages’ describes a setting with an opponent that causes random outages present in the OOD dataset.
The agents can be ordered by performance: Do-Nothing $<$ Naive $<$ Verify $<$ Greedy Hybrid $<$ N-1 Hybrid. The naive agents (rows 4 to 6) and verify agents (rows 8 to 10) come close to the performance of the N-1 expert agents (row 3). The hybrid agents (rows 12 to 14, and 16 to 18) can match the N-1 expert agents.
On the setting with no or ID outages (columns ’Full Network’ and ’ID Outages’) the HetGNN network consistently performs best, followed by the FCNN and, finally, the HomGNN (rows 4 to 6, and 8 to 10). Additional simulation reduces or nullifies the differences (rows 12 to 14, and 16 to 18).
On the setting with OOD outages (column ’OOD Outages’), the GNNs outperform the FCNN (rows 4 to 6, and 8 to 10). However, additional simulation also decreases this effect (rows 12 to 14, and 16 to 18). Furthermore, the performance of the models trained on the ID dataset remains substantially lower than models trained on OOD dataset (rows 7 and 11), even with additional simulation (rows 15 and 19).
5.6 Efficiency
Finally, we consider the inference speed of the different agents. The inference speed was measured on an Apple M1 CPU with minimal background processes on the first fifty validation scenarios. Table 5.6 lists the durations per model type and agent. Figure 5.6 plots these values against the agent’s performance on the full-network topology. The GNNs, particularly the HetGNNs, are considerably slower than the FCNNs. However, these differences are small relative to the simulation times of the expert agents, which are orders of magnitude higher. Most importantly, all GNN agents have a favorable combination of speed and performance.
@@ 5.6 Efficiency
tableInference durations of various agents. The hybrid models all use the FCNN. The differences between model types for the hybrid agents are relatively insignificant. Agent Duration ($\upmu$s) Greedy 6.77E5 N-1 4.70E6 Naive FCNN 5.74E2 Naive HomGNN 2.81E3 Naive HetGNN 3.84E3 Verify FCNN 1.14E4 Verify HomGNN 1.50E4 Verify HetGNN 1.58E4
@@ 6 Discussion
The heterogeneous GNN consistently performs best, followed by the FCNN, and lastly, the homogeneous GNN. This is true for both the accuracy and grid operation. The superiority of the HetGNN to the FCNN indicates that GNNs can exploit the graph structure better. The inferiority of the HomGNN can be attributed to the busbar asymmetry problem. However, smoother graphs and longer path lengths might also play a role. The HetGNN’s coarser graphs can be attributed to the more diverse edge representation. The GNNs take considerably longer to evaluate than the FCNNs, although this difference is very tiny compared to the expert agents.
The GNNs achieve far higher accuracy than the FCNNs on the OOD networks. This is also reflected in the simulation performance, although the effect is smaller. Although GNNs show a superior ability to generalize to OOD networks, they still perform considerably worse than GNNs trained on these networks.
@@
There seemed to be a limit to the accuracy of the trained models. This can be attributed to class imbalance and class overlap. We hypothesize that the class overlap originates from the Grid2Op simulate function. It is possible that identical network states have diverging forecasts and consequently diverging actions. The limited accuracy does not stop the ML models from successfully completing many of the grid operation days. The naive ML models come close to the performance of the expert agents. The hybrid agents can match them.
Hybrid agents can outperform expert agents in specific settings, which is surprising. We hypothesize that this results from omitting the state-action pairs from failed runs. The model bias towards frequent actions might also result in more robust topologies.
An interesting effect is that model differences are more pronounced in accuracy than simulation performance. Similarly, the limited accuracy does not stop the models from operating the grid well. We have two explanations for this. Firstly, a misprediction does not necessarily lead to a game-over. This is probably particularly true in regions with class overlap. Secondly, difficult days led to the generation of many relatively uncommon datapoints. Failure to predict many such datapoints would consequently affect the completion of only few days.
7 Future Work
Our findings indicate the promise of fully graphical neural networks for topology control methods. However, further research is necessary to establish this. Future work should compare fully GNN models with models that only use GNNs for feature extraction. Similarly, the heterogeneous graph representation should be introduced into other approaches. Moreso, it would be good to apply these methods in a reinforcement learning framework.
Further efforts are likely also necessary to scale this approach to larger grids. Fully GNN models might require very deep GNNs to model large graphs. Very deep GNNs are associated with additional challenges [46]. Future research could focus on methods to downsample grid graph representations, e.g., through trainable graph pooling [12]. Alternatively, it might be beneficial to combine GNN layers and other layers to capture long-range patterns.
The imitation learning approach used here can also be developed further. Our approach suffers from distribution shift, a compounding divergence between the behavior of the ML and the expert agent [1]. More advanced IL frameworks, such as DAgger, address this [34]. Investigating other expert agents and developing an improved dataset would also be valuable.
Finally, the generalization capabilities of graph neural networks should be studied more extensively. Future studies should investigate how factors such as the diversity of training data and differences between the ID and OOD data affect generalization.

### v3

@@ 2.4 Regimes
To evaluate performance under varying operational stress, three environmental regimes were considered. The full-network regime does not involve line outages. The planned-outage regime incorporates static line outages. The unplanned-outage regime simulates stochastic line outages using the opponent described by Manczak et al. manczak2023hierarchical . This opponent disables a randomly selected line for a four-hour duration, twice daily, with a minimum one-hour interval between outages.
@@ 3.1 Datasets
We use the dataset generated by the greedy and N-1 expert agents, described in our previous work deJong2024 . Both agents use extensive load flow simulations to find the optimal action. The greedy agent selects actions by minimizing the maximum line loading of the grid in its current state, whereas the N-1 agent minimizes the maximum line loading under the disablement of any power line. The N-1 agent demonstrated superior performance compared to the greedy agent in settings both with and without outages. Both agents use an activity threshold below which “do-nothing” actions are automatically selected. State-action pairs from such do-nothing actions, or from episodes that the expert agents could not successfully complete, were excluded from the dataset. Do-nothing actions selected above the activity threshold were included.
We constructed two datasets. First, an in-distribution (ID) dataset, containing network configurations on which the models are trained and evaluated. This dataset was used in our previous study deJong2024 . This dataset incorporates state-action pairs from both the greedy agent and the N-1 agent. Specifically, it incorporates state-action pairs from the N-1 agent operating on the full network, and from the greedy agent operating on a subset of N-1 networks (lines 0, 2, 4, 5, 6, and 12 disabled). These specific N-1 networks were selected because the greedy agent can operate them well deJong2024 .
@@
For both datasets, we split the datapoints into 70/10/20 train/validation/test sets based on scenario. The partition sizes are listed in Table 2. In this paper, when we refer to “ID networks” or “OOD networks”, we refer to the network variations in the ID or OOD datasets, respectively. Likewise, when we refer to “ID outages” or “OOD outages”, we refer to the outages present in respectively ID networks or OOD networks.
@@
Table 2: The number of datapoints in the datasets and train/validation/test sets.
@@ 3.2 Datapoints
Each datapoint includes each object’s features, the topology vector, and the expert action. The features per object type are listed in Table 3. Object features are normalized. If a line is disabled, its endpoint features are zero-imputed, and the corresponding topology vector values are -1. Actions are converted from a set-action format into a switch-action format. Thus, actions are represented by a vector with a length of the topology vector. Each index indicates whether the corresponding object is switched between busbars (a value of 1) or not (a value of 0). This presents a multi-label binary classification task.
@@ 3.3 FCNN
The fully connected neural network (FCNN) model consists of an input layer, multiple hidden layers, and an output layer. The input vector is constructed by flattening the object features into a single vector and concatenating the topology vector. The hidden layers use the ReLU activation function. The output layer uses the sigmoid activation function to constrain the output to the $(0,1)$ range. The output vector has the length of the topology vector, so one value $p_{u}\in(0,1)$ is predicted per node $u$.
@@ 3.4 Homogeneous GNN
Applying a graph neural network to the power grid requires a graph representation of the grid. We represent the grid objects as nodes within the graph. Edges are set between nodes if the corresponding objects are connected to the same busbar or represent (corresponding) line endpoints. We call this graph representation homogeneous as it does not consider edge types. We call the associated GNN the HomGNN. Figures 2a and 2b respectively display an example grid and its homogeneous graph representation.
@@
[…]
where $\textbf{h}_{u,0}$ and $\textbf{x}_{u}$ are, respectively, the initial node embedding and the features of node $u$. The embeddings of subsequent layers are computed with the message passing rule 4700287 :
@@ 3.5 Busbar Information Asymmetry
The aforementioned graph neural network formulation suffers from a problem. The output of each node specifies whether to switch the corresponding object across busbars. However, messages from objects on the current busbar are passed, while messages from the objects on the other busbar are not. We call this the busbar information asymmetry. We posit that information about objects on both busbars should be equally available, as both groups of objects are involved in actions. Moreover, this decreases the model’s expressiveness: objects at the same substation but on different busbars and objects at entirely different substations cannot be distinguished. Lastly, the absence of inter-busbar connections also decreases the graph’s connectivity. This can result in longer paths that may require deeper GNNs.
@@ 3.6 Heterogeneous GNN
To address this problem, we propose a heterogeneous graph representation, and a corresponding GNN, the HetGNN This heterogeneous representation features three edge types: edges connecting objects on the same busbar, edges connecting objects on the other busbar but within same substation, and edges connecting corresponding line endpoints. The heterogeneous graph representation is displayed in Figure 2c. The message passing rule is adapted to consider a weight and neighborhood aggregation per edge type:
@@
where $\mathcal{N}_{same}$ and $\mathcal{N}_{other}$ denote the neighborhoods of objects on the same and other busbar, respectively. The $W_{line,k}\textbf{h}_{line,k}$ term represents the addition of the embedding of a node connected by a power line. This term is therefore omitted for nodes that do not represent line endpoints. Figure 2e displays an example of heterogeneous message passing.
@@ 3.7 Optimization & Postprocessing
where $\textbf{w},\textbf{y}$, and p denote the label weight, target, and prediction vectors, respectively. We introduce label weights to prevent the prediction vectors from collapsing to zeros. This occurs as the majority of target values are zero. A lower label weight $0<\alpha<1$ is assigned to labels that do not correspond to objects at either the target or predicted substation:
@@
There is no target substation if the target action is a do-nothing action. The predicted action is classified as a do-nothing action, without a predicted substation, if all predictions $p_{i}\in\textbf{p}$ do not exceed $0.5$. Otherwise, the predicted substation is the substation where the predictions $\textbf{p}_{s}$ at that substation $s\in S$ maximize
@@ 3.8 Hyperparameter Tuning & Training
Figure 3: The training curves of the five models per model type.
@@ 3.9 Imitation Learning Agents
The ML models are applied to the environment. They can be applied directly or combined with simulation. The naive agent executes a model’s predicted action directly. We observe that this agent occasionally fails by predicting a singular erroneous action. The verify agent addresses this by verifying predicted actions with simulation. The predicted action is simulated. A do-nothing action is selected if the simulated line loadings are increased beyond the thermal limit. We also consider hybrid agents. The verify+greedy agent normally functions as the verify agent, but switches to the greedy agent if a line breaches the thermal limit. The verify+N-1 agent functions similarly but with the N-1 agent. Each agent also uses the activity threshold used by the expert agents deJong2024 .
@@ 4.1 Supervised Learning
Table 5 presents the accuracies of the different models across the data partitions and the ID/OOD network groups. Figure 4 displays the test accuracies of the models on both datasets. On the ID dataset (rows 1-3), the accuracies remain limited to approximately 80%. On each split, the HetGNN achieves the highest accuracy, followed by the FCNN and, lastly, the HomGNN. A minor decrease in accuracy is observed between the training, validation, and testing sets for all model types, suggesting a slight degree of overfitting. As shown in row 4, accuracy decreases marginally when the post-processing step outlined in Section 3.7 is omitted.
The default topology, characterized by all objects connected to the same busbar, appears frequently within the ID dataset, constituting 46% of the test set. This configuration avoids the busbar asymmetry problem, thereby negating the theoretical advantage of the HetGNN over the HomGNN. As shown in row 5, the accuracies on the standard topology are relatively similar. Conversely, as shown in row 6, the accuracies vary considerably in topologies with split busbars.
@@
The GNNs achieve far higher accuracies on the OOD dataset than the FCNNs (rows 7-9). The HetGNNs also achieve higher accuracies than the HomGNNs. However, the performance of either the HomGNN or the HetGNN on the OOD dataset is still substantially lower than models specifically trained on OOD data (column ’OOD-GNN’).
@@ 4.2 Error Analysis
We repeated the error analysis previously performed on the FCNN deJong2024 for both GNN model types. As with the FCNN, class imbalance and overlap contribute to the limited accuracy. As illustrated in Figure 5, both GNN types exhibit a tendency to under-predict infrequent classes. We investigated frequently confused class pairs, i.e. classes commonly mistaken for each other during prediction. Figure 6 indicates that the errors associated with these frequently confused classes occur within overlapping regions of the feature space.
@@
Analysis of nearest neighbors further supports the hypothesis that class overlap contributes to reduced accuracy. We applied one of the HetGNNs to a subset of 2,500 validation datapoints. For data points whose nearest neighbor belongs to the same class, the model achieves an accuracy of 92.97%. In contrast, for data points whose nearest neighbor belongs to a different class, the accuracy drops to 44.44%. Similar trends were observed across models.
@@ 4.3 Graph Smoothness
Given the well-documented tendency of graph neural networks to exhibit oversmoothing chen_oversmoothing , 10.5555/3504035.3504468 , which can degrade model performance, we assess smoothness of the graphs produced by both GNNs. We calculate the mean average distance (MAD) values, which quantify the distance between neighboring node embeddings. Lower MAD values correspond to higher embedding similarity and thus increased graph smoothness. Figure 4.4 shows the MAD values over the various layers for both GNN types, calculated on the first thousand validation datapoints. As shown, the MAD values of the middle GNN layers are consistently higher in the HetGNN than in the HomGNN. This finding is consistent across the different models per type, suggesting that oversmoothing is a smaller concern for the HetGNN.
@@ 4.4 Graph Diameter
We mentioned that the heterogeneous graph representation can lead to shorter paths. We measure this by the diameter, i.e., the shortest path length between the two most distant nodes. We calculate the diameter by finding the lowest exponent of the adjacency matrix that produces a matrix without zero entries Duncan2004PowersOT . The diameters of the 25 most common topologies are shown in Figure 4.4. As visible, the heterogeneous graph representation infrequently leads to a shorter diameter.
@@ 4.5 Simulation Performance
Table 6 shows the percentage of days completed per agents across the different settings. The ‘Full Network’ column describes the setting without outages. The ‘ID Outages’ and ‘OOD Outages’ columns correspond to settings with random outages derived from the in-distribution and out-of-distribution datasets, respectively (see Sec. 3.1).
Agent performance follows a consistent order: Do-Nothing << Naive << Verify << Greedy Hybrid << N-1 Hybrid. The naive agents (rows 4-6) and verify agents (rows 8-10) achieve performance near to the N-1 expert agent (row 3). The hybrid agents (rows 12-14 and 16-18) can match the performance of the N-1 expert agent.
On the setting with no or ID outages (columns ’Full Network’ and ’ID Outages’) the HetGNN network consistently performs best, followed by the FCNN and, lastly, the HomGNN (rows 4-6 and 8-10). Additional simulation reduces or nullifies the differences (rows 12-14 and 16-18).
On the setting with OOD outages (column ’OOD Outages’), the GNNs outperform the FCNN (rows 4-6 and 8-10). However, additional simulation also reduces this effect (rows 12-14 and 16-18). Furthermore, the performance of the models trained on the ID dataset remains substantially lower than models directly trained on OOD dataset (rows 7 and 11), even with additional simulation (rows 15 and 19).
4.6 Computational Efficiency
We finally consider the inference speed of the different agents. Inference times are measured using an Apple M1 CPU with minimal background processes across the first fifty validation scenarios. Table 4.6 lists the mean duration per model and agent type. Figure 4.6 plots these values against the agent’s performance on the setting with OOD outages. The GNNs, particularly the HetGNNs, are considerably slower than the FCNNs. However, their computation times remain orders of magnitudes smaller than those of the expert agents. Most importantly, all GNN agents have a favorable combination of speed and performance.
@@ 4.6 Computational Efficiency
[…]
@@ 5.1 Interpretation
The heterogeneous GNN consistently achieves the highest performance, in regards to both classification accuracy and grid operation ability. The homogeneous GNN exhibits the lowest performance, being surpassed by the FCNN. The superior performance of the HetGNN compared to the FCNN demonstrates the ability of GNNs to effectively exploit graph structure. The inferior performance of the HomGNN can be attributed to the busbar asymmetry problem. Smoother graphs and longer path lengths might further contribute the inferior performance. The more diverse edge representation of the HetGNN likely reduces the oversmoothing effect of GNNs. Concerning efficiency, the GNNs take considerably longer to evaluate than the FCNNs. However, this difference is negligible in comparison with the expert agents.
A question that naturally arises is why homogeneous graph neural networks have exhibited good performance in related research yoon2020winning . We believe that the answer lies in the fact that models used in other research combine GNN layers with other layers (e.g., linear layers), which mitigates the busbar asymmetry problem. This presents an interesting topic for future work.
GNNs demonstrate a superior ability to generalize to out-of-distribution (OOD) network configurations than FCNNs. This effect is present in both accuracy and, albeit with reduced magnitude, simulation performance. Despite their superior generalizability, GNN OOD performance remains substantially below that of models directly trained on OOD network configurations.
@@
The accuracy of the trained models appears to reach a plateau, which may be attributed to issues such as class imbalance and class overlap. We hypothesize that the observed class overlap originates from the Grid2Op simulate function. It is possible that identical network states have diverging forecasts and consequently diverging actions, causing class overlap. This limitation in classification accuracy does not preclude the models from achieving good operation performance, however. The naive ML agents achieve operational performance close to the expert agents, whereas the hybrid agents can match them.
Counterintuitively, hybrid agents sometimes achieve performance superior to the expert agents. A potential source for this effect is the exclusion of samples from the dataset for days that the expert agents could not successfully complete, positively biasing the models to successful behavior. The model bias towards frequent actions might also result in robust topologies that can accommodate a wider set of injection profiles.
An interesting effect is that model differences are more pronounced with respect to accuracy than operational performance. Similarly, the limited accuracy does not stop the models from effectively operating the grid. We propose two contributing factors. First, a misprediction does not necessarily lead to a game-over; the class overlap suggests that there are often multiple viable actions. Second, certain challenging days generate a high volume of uncommon data points. Misclassification of these datapoints affects a considerable subset of the dataset but is limited to only few operational days.
5.2 Future Work
We propose the integration of the methods developed here, particularly the heterogeneous graph representation, into other approaches as a focus of future research. The heterogeneous graph representation could be integrated seamlessly into previously proposed model architectures. It is particularly still an open question how HetGNN layers interact with non-graphical model layers, warranting investigation. Incorporating the method into a reinforcement learning framework is also a crucial next step.
It is particularly promising to integrate these approaches into a multi-agent reinforcement learning (MARL) framework, which has recently gained attention for topology control de_Mol_2025 , vandersar2023 . Within MARL, the combinatorially large action space can be factorized across sub-agents. The agents developed here could be deployed as localized sub-agents within such a system, offering a natural path toward scaling to larger grids.
Although existing approaches have applied imitation learning only in it’s simplest form, there are compelling reasons to apply more advanced methods. Notably, the simple behavioral cloning learning setup suffers from distribution shift, a compounding divergence between the behavior of the ML and expert agent belkhale2024data . DAgger is a more advanced method that addresses this problem pmlr-v15-ross11a . Additionally, it would also be valuable to further explore expert agents and develop better datasets.
Finally, given the dynamic nature of power grids, generalization of machine learning models remains a topic to be studied more extensively. There is yet little knowledge about how factors such as the diversity of training data and differences between ID and OOD data affect generalization. Data augmentation might be a promising technique to obtain models with wide applicability.
6 Conclusion
To guarantee energy security, transmission system operators (TSOs) must effectively manage grid congestion. Topology control guided by machine learning presents a powerful tool for TSOs to do so. While graph neural networks (GNNs) are well-suited to model the grid’s structure, we identify that their straightforward application suffers from a critical limitation: the busbar asymmetry problem. To address this, we introduce a heterogeneous graph representation that resolves the asymmetry and enables effective learning. Heterogeneous graph neural networks outperform homogeneous graph neural network, the latter of which fails to outperform the fully-connected neural network (FCNN) baseline. In turn, heterogeneous GNNs also surpass fully connected neural networks (FCNNs), demonstrating the value of incorporating grid topology explicitly. This performance gap is most pronounced in out-of-distribution settings, where FCNN performance deteriorates significantly, while GNNs retain a higher degree of generalization. Importantly, agents trained with these methods achieve operational performance comparable to expert systems, while offering greatly improved computational efficiency. These findings demonstrate the potential of GNN-based agents as a solution to support TSOs in real-time grid management.